OpenStudy (anonymous):

Tests by an independent auditing firm show that 70% of students in a certain school think that Algebra is the best course they have ever taken. Use the entire line of random numbers below to estimate the probability that a randomly selected group of five students will contain at least 4 Algebra lovers. Let the digits 1-7 represent a student who thinks that Algebra is the best course. 12024 55976 61475 70726 25408 62279 71874 03499 92659 26041 Convert your probability answer to a whole number representing a percent.

OpenStudy (anonymous):

So you're trying to find the probability that a group of numbers will contain at least four numbers between 1 and 7: 12024 Yes 55976 Yes 61475 Yes 70726 Yes 25408 Yes 62279 Yes 71874 Yes 03499 No 92659 No 26041 Yes So 8 out of the 10 groups satisfy this condition. 8/10=4/5 chance

OpenStudy (anonymous):

Oh you wanted percent... 4/5=80%

OpenStudy (anonymous):

Thank you!! I don't want to sound rude, but I need to know FOR SURE this answer and someone else told me it was 50%. Can anyone else help with this question just to be sure? THANKS!

OpenStudy (anonymous):

Oh no problem, I've made stupid mistakes in the past so I won't doubt it ^^

OpenStudy (anonymous):

@skinny23 @luisloco25 @amistre64 @jim_thompson5910 @nathan917 @numb3r1

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (anonymous):

@amistre64 @amistre64 @amistre64 I need you!

OpenStudy (amistre64):

the mean of the means

OpenStudy (amistre64):

lets trim the number, we know each sample is size is 5, to lets see what the ratio are for 1-7 1224 5576 61475 7726 254 6227 7174 34 265 2641 4/5 4/5 5/5 4/5 3/5 4/5 4/5 2/5 3/5 4/5 what is the mean of these proportions?

OpenStudy (amistre64):

the mean proportion of the samples is 74/100

OpenStudy (anonymous):

Oh right, I forgot 0's don't count ^^ Good thing you got a smart person in here to help =P

OpenStudy (amistre64):

contain at least 4 Algebra lovers .... yeah, i didnt see that at first :) so out of the 10 samples, we have 7 that 4 or 5 7/10 = 70%

OpenStudy (amistre64):

im not sure if there is more to this, like making a confidence interval, and comparing the sample mean with the other stuff .....

OpenStudy (anonymous):

Let's not go there... This seems like strictly Algebra ^^

OpenStudy (amistre64):

:) no field and ring theory? darn

OpenStudy (anonymous):

And I was getting so excited too! (sarcasm)

OpenStudy (anonymous):

70%